OpenStudy (anonymous):

2m-nx=x+4, solve for x

OpenStudy (anonymous):

solve for x means that we want x on one side lets have it on the right side. that means we first add nx to both sides and end up with 2m=x+4+nx now we subtract 4 to both sides and get 2m-4=x+nx This step is confusing but on the right side we must distribute out an x see if you can try that .